It takes a small pump 5 minutes longer than a large pump to pump all of the blood out of an average turnip. If both pumps are hooked to the turnip, they can do the job in only 6 minutes. How long does it take each pump working by itself?
Answer by richwmiller(17219)

6/x+6/y=1,
y=x+5
x = 10, y = 15
6/15=2/5
6/10=3/5